Find the eigenvalues and eigenfunctions of the following Sturm-Liouville probelm (e^(2x)y')' + e^(2x) (λ + 1)y = 0; y(0) = 0 = y(π). 


1
Expert's answer
2021-10-05T17:06:06-0400

(e2xy)+e2x(λ+1)y=0e2xy+2e2xy+e2x(λ+1)y=0Divide through by e2xy+2y+(λ+1)=0We can only have a solution when λ>0m2+2m+(λ+1)=0m=1±λy(x)=e1[C1cos(xλ)+C2sin(xλ)]Apply the boundary conditionsy(0)=e1C1    C1=0y(π)=e1[C2sin(πλ))C20    sin(πλ)=0    πλ=nπ   n=1,2,3,    λn=n2This is the Eigen valuesyn(x)=e1[Cnsin(nx)]This is the Eigen Functions(e^{2x}y')'+e^{2x}(\lambda +1)y=0\\ e^{2x}y''+2e^{2x}y''+e^{2x}(\lambda +1)y=0\\ \text{Divide through by } e^{2x}\\ y''+2y'+(\lambda + 1)=0\\ \text{We can only have a solution when } \lambda >0\\ m^2+2m+(\lambda + 1)=0\\ m=-1\pm\sqrt{-\lambda}\\ y(x)=e^{-1}[C_1\cos(x\sqrt{\lambda})+C_2\sin(x\sqrt{\lambda})]\\ \text{Apply the boundary conditions}\\ y(0)=e^{-1}C_1 \implies C_1=0\\ y(\pi)=e^{-1}[C_2\sin(\pi \sqrt{\lambda}))\\ C_2 \neq 0\\ \implies \sin(\pi \sqrt{\lambda})=0\\ \implies \pi \sqrt{\lambda}=n\pi ~~~n=1,2,3, \cdots\\ \implies \lambda_n =n^2 \text{This is the Eigen values}\\ y_n(x)=e^{-1}[C_n\sin(nx)]\\ \text{This is the Eigen Functions}